OpenStudy (anonymous):

The area of a rectangular room is given by the trinomial x^2 - 3x - 28. What are the possible dimensions of the rectangle? Use Factoring. A.(x-6) and (x-5) B.(x+6) and (x-5) C.(x+6) and (x+5) D.(x-6) and (x+5)

OpenStudy (jtvatsim):

Something about the question doesn't make sense... are you sure those are the answer choices?

OpenStudy (jtvatsim):

Because there doesn't seem to be any way to factor x^2 - 3x - 28 into the given choices.

OpenStudy (anonymous):

A.(x+7) and (x-4) B.(x-7) and (x-4) C.(x+7) and (x+4) D.(x-7) and (x+4) these are the correct answer choices @jtvatsim

OpenStudy (jtvatsim):

OK, that makes much more sense.

OpenStudy (anonymous):

Can you help?

OpenStudy (jtvatsim):

So how would you start to factor x^2 - 3x - 28? Any ideas?

OpenStudy (anonymous):

No, I actually needed help with the entire problem.

OpenStudy (jtvatsim):

OK... Let's start with something easier and work our way up. I will give you a different question to answer first. Are you able to multiply these two things together: x(x + 1) = ?

OpenStudy (anonymous):

1x

OpenStudy (jtvatsim):

OK, and how did you get that?

OpenStudy (anonymous):

x*x =x x+ 1+ 1x

OpenStudy (anonymous):

x+1 =1x *

OpenStudy (jtvatsim):

OK, I see what you did. But, we have to remember that x*x = x^2. The exponent tells us how many times we multiplied x., so It's important to keep that there.

OpenStudy (jtvatsim):

So here is what happens: x(x+1) = x*x + x*1 Notice how the left x multiplies both pieces inside. Then, x*x becomes x^2, and x*1 becomes x. So the final answer is x^2 + x. x(x+1) = x^2 + x. Does that make sense?

OpenStudy (anonymous):

yes, it makes sense

OpenStudy (jtvatsim):

OK, how about this one just to make sure you've got it: x(x + 3) = ? What would you get for that?

OpenStudy (jtvatsim):

I know this is painful, but it will help us solve the real question we care about. :)

OpenStudy (anonymous):

x^2+x3

OpenStudy (anonymous):

3x